OpenStudy (anonymous):

Rational numbers word problems -- Problems to be typed in reply box; there are 10.

OpenStudy (anonymous):

hi. the first problem is

OpenStudy (anonymous):

jane is baking a cake. the bottom layer is 3/4 of an ich, the two middle layers are 1/2 of an inch each, and the top layer is 1 5/6 inch

OpenStudy (anonymous):

how tall is the cake.

OpenStudy (anonymous):

Please help.

OpenStudy (tyteen4a03):

Please ask a separate question for each question. The expression you're looking for is \(\frac {3}{4} + 2\frac{1}{2} + \frac{11}{6}\). Simplify 2*(1/2) and it becomes 1. Now find the GCD of 4 and 6.